OpenStudy (anonymous):

@Arnab09 Factor 8x^2 + 2x – 3. A. (2x – 1)(4x + 3) B. (2x + 3)(4x – 1) C. (8x + 1)(x – 3) D. The polynomial is prime

OpenStudy (anonymous):

8x^2+6x-4x-3= 2x(4x+3)-2(4x+3)=2(x-1)(4x+3)

OpenStudy (anonymous):

that would be A.?

OpenStudy (anonymous):

sorry, typing mistake there.. 2x(4x+3)-1(4x+3)=(2x-1)(4x+3)

OpenStudy (anonymous):

can u help me with one more?

OpenStudy (anonymous):

post another question

OpenStudy (anonymous):

ok